OpenStudy (anonymous):

A manufacturer produces bolts of a fabric with a fixed width. The quantity of this fabric (measured in yards) that is sold is a function of the selling price p (in dollars per yard), so we can write q = f(p). Then the total revenue earned with selling price p is R(p) = pf(p). a. What does it mean to say that f(20) = 10,000 and f'(20) = -350? b. Assuming the values in part (a), find R'(20) and interpret your result

OpenStudy (anonymous):

@zepdrix can you please help me with this question?

zepdrix (zepdrix):

Hmm interesting problem :) Thinking..

OpenStudy (anonymous):

oh okay... good luck....

zepdrix (zepdrix):

Let's try to interpret the f(20) first. Recall that this function is defined to be our `quantity`. And the function depends on the `price`. So at a price of p=20, we are able to sell 10000 bolts of fabric. f(20)=10000.

OpenStudy (anonymous):

okay.... I get that now

zepdrix (zepdrix):

The f'(20) is a little trickier. The derivative represents the instantaneous rate of change of the function, the quantity function, at p=20. So at a price of p=20, The amount of bolts of fabric we are selling is decreasing by 350 bolts per... dollar maybe?

OpenStudy (anonymous):

That should be it

OpenStudy (anonymous):

it makes a lot of sense to put it that way

zepdrix (zepdrix):

Oh they gave us the units at the start :) I forgot about that. So our at a price of $20, Our quantity is decreasing by 350 yards of fabric per dollar. Ya ya ya, I think that makes sense.

zepdrix (zepdrix):

For part b, \[\large\rm R(p)=p~f(p)\]To get the derivative of the revenue function, you'll need to apply your product rule. Remember that one? :)

OpenStudy (anonymous):

I understand that.... i really don't know how to start

OpenStudy (anonymous):

p is not a function, it is a constant. is it necessary to use the product rule?

zepdrix (zepdrix):

I guess it's not a constant, p is price. Price can range from any number of values that we set it at. So price is a variable. It `will be` constant, when we set it at $20. But we have to apply our differentiation rule `before` we do that.

zepdrix (zepdrix):

What we're doing is, looking for a derivative function `with respect to p`.\[\large\rm R'=\frac{dR}{dp}\]

OpenStudy (anonymous):

oh okay.... i get it

zepdrix (zepdrix):

Applying Product Rule gives us this "set up",\[\large\rm R'(p)=p'~f(p)+p~f'(p)\]That rule looks familiar, yes? :) and hopefully you can make sense of the way p' will simplify. It's just the derivative of p with respect to p. Similar to if you had taken a derivative like this,\[\large\rm \frac{d}{dx}x=1\]

zepdrix (zepdrix):

So our Derivative Revenue Function looks like this,\[\large\rm R'(p)=1\cdot f(p)+p~f'(p)\]\[\large\rm R'(p)=f(p)+p~f'(p)\]

zepdrix (zepdrix):

\[\large\rm R'(\color{orangered}{p})=f(\color{orangered}{p})+\color{orangered}{p}~f'(\color{orangered}{p})\]And we want to evaluate this derivative at p=20,\[\large\rm R'(\color{orangered}{20})=f(\color{orangered}{20})+\color{orangered}{20}~f'(\color{orangered}{20})\]

zepdrix (zepdrix):

To do so, you'll need to use the `pieces of information they gave you` from before:\[\large\rm f(20)=10000,\qquad\qquad f'(20)=-350\]

OpenStudy (anonymous):

Thank you very much I get it now I got my answer to be 3000 dollars

zepdrix (zepdrix):

Yay good job \c:/ np